# A predictive model for bubble–particle collisions in turbulence

### v1

@@ 2 Modelling the bubble slip velocity probability density function
[…]
@@
We note that numerically evaluating (13) directly at large $\langle w_{b}\rangle$ may lead to undefined values as $\lim_{x\to\infty}\sinh{(x)}/x\to\infty$ and $\lim_{x\to\infty}\exp{(-x^{2})}\to 0$. This corresponds to the strong gravity regime where the slip velocity vector is dominated by its vertical component, such that $f(w_{b})$ converges to a normal distribution centred at $\langle w_{b}\rangle$ (see C). Therefore, we use
$f(w_{b})$ =$=\begin{cases}\,\frac{\sqrt{2}w_{b}}{\sqrt{\pi}\sigma\langle w_{b}\rangle}\exp{\bigg{(}-\frac{w_{b}^{2}+\langle w_{b}\rangle^{2}}{2\sigma^{2}}\bigg{)}}\sinh{\bigg{(}\frac{w_{b}\langle w_{b}\rangle}{\sigma^{2}}\bigg{)}}&\text{if $\langle w_{b}\rangle/\sigma\leq 16$,}\\ \,\frac{1}{\sqrt{2\pi}\sigma}\exp\bigg{(}-\frac{(w_{b}-\langle w_{b}\rangle)^{2}}{2\sigma^{2}}\bigg{)}&\text{otherwise.}\end{cases}$ (14)
@@
2.1 Mean vertical slip velocity
@@
$\langle w_{b}\rangle=\left\{\begin{aligned} v_{q}(1-\chi Fr_{L}^{2})\quad&\text{for $1/Fr_{L}\geq 2.08$,}\\ 0.37\frac{v_{q}}{Fr_{L}}\quad&\text{otherwise,}\\ \end{aligned}\right.$ (16)
@@
2.2 Standard deviation of the slip velocity
@@
where the response function
@@
Note that as the expression of $f_{b}$ used tacitly implies $Re_{b}=2r_{b}\sigma_{i}/\nu$, so (23) should only apply for $01/Fr=$ 0.
@@
where $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ (Nguyen and Schulze, 2004) with $\langle w_{b}\rangle$ defined by (16).
@@ 3.1 Equations and solvers
$\langle\mathbf{f}^{ibm}\rangle_{\Pi}(t)=\frac{1}{||\Pi||}\int_{\Pi}\mathbf{f}^{ibm}(\mathbf{x},~{}t)\rm{d}\mathbf{x}$ (33)
at every time step, where $\Pi$ denotes the entire computational domain, $\langle\cdots\rangle_{\Pi}$ indicates spatial averaging over $\Pi$, and $||\Pi||$ is the volume of $\Pi$. Hence, the force exerted by the bubble on the fluid is
@@
In contrast to the bubbles, the size of the mineral particles in the flotation processes are comparable to the Kolmogorov length scale $\eta$, meaning they can be modelled as point-like particles. Here, we consider one-way coupled non-settling particles that can overlap with each other and only experience the drag, pressure gradient, and added mass forces, which is consistent with the modelling assumptions in Jiang and Krug (2025). The particle dynamics is therefore governed by (15), except all bubble quantities are replaced by those of the particles, the gravity term is ignored, and $f_{p}=1+0.15Re_{p}^{0.687}$. The particle equation is solved using a finite-difference scheme where the spatial derivatives are discretised using a second-order central finite-difference method and time marching is performed with a second-order Adams–Bashforth scheme using the same time step size as the fluid solver. Since the terms involving $\mathbf{u}$ need to be evaluated at particle positions, their values are interpolated from the Eulerian grid using a tri-linear scheme.
@@ 3.2 Simulation parameters
The equations described in §3.1 are solved to simulate bubbles with $\rho_{b}/\rho_{f}=1/1000$, $St_{b}=0.5-6.3$ and $101/Fr=1-$ 10 in homogeneous isotropic turbulence with $Re_{\lambda}=64$. Table 1 lists the parameters of the simulations and the turbulence statistics. We limit ourselves to low bubble volume fractions $\phi$ ($\leq 10$ bubbles) where bubbles rarely approach each other. Nonetheless, in case this happens, we employ a soft collision model with a coefficient of restitution of 0.8 to properly capture the lubrication force at small separations and prevent coalescence (Costa et al., 2015). Our simulations are performed in a triply periodic domain. Hence, in principle it is possible for a rising bubble to encounter its own wake, which would affect the statistics. In the wake, the velocity defect $\delta u_{b}$ decays initially with the downstream distance $z$ as $z^{-1}$ until it reaches the same order of magnitude as the turbulent velocity fluctuations at $z_{0}$, beyond which it decays as $z^{-2}$ (Legendre et al., 2006). We take the velocity defect at the bubble position $\delta u_{b}(0)=\langle v_{b,z}\rangle$, define $z_{0}$ by $\delta u_{b}(z_{0})=u^{\prime}$, and estimate the velocity defect one domain length downstream of the bubble $\delta u_{b}(L_{z})=0.01u^{\prime}\ll u^{\prime}$, which is negligibly small. Additionally, we ran an extra simulation with $(St_{b},1/Fr)=(6,10)$ in a longer domain and find that the statistics do not change significantly. These considerations suggest that the bubble’s motion is not significantly affected by its own wake. To ensure sufficient spatial resolution of the boundary layer on the bubbles, we estimate the thickness of the boundary layer at the front stagnation point as $\delta/(2r_{b})=1.13/Re_{b}^{1/2}$ following Johnson and Patel (1999), and resolve this by at least 7 grid points. A grid convergence study (see D) confirmed that this is sufficient for the present cases. The bubbles are initialised on a primitive cubic lattice with the fluid velocity at their positions and allowed to adapt to the flow for 10 large eddy turnover times before their statistics are collected. The instantaneous fluid velocity at bubble position is determined by averaging the fluid velocity within a distance of $3r_{b}$ from the bubble, which for the case of a bubble subjected to a steady uniform flow recovers a value that reaches 90% of the ground truth (Kidanemariam et al., 2013).
@@ 4.1 Mean vertical bubble slip velocity
We compare the predicted values of the mean vertical bubble slip velocity $\langle w_{b}\rangle$ with the actual values in figure 2(a). The data shows that $\langle w_{b}\rangle$ increases with $St_{b}$ and $1/Fr$, which is nicely captured by our model qualitatively and quantitatively. This justifies our use of (16) to model $\langle w_{b}\rangle$, even though (16) was developed for the bubble rise velocity instead of the slip velocity in turbulence. We additionally review the mechanism that reduces the bubble rise velocity in turbulence for the different cases by showing the values of $1/Fr_{L}$ in figure 2(b). For all the tested parameters except $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,10)$, $1/Fr_{L}$ is less than the transition value of 2.08 that separates the two scalings in (16). This suggests that for the vast majority of the simulated cases, nonlinear drag plays a dominant role in reducing the mean vertical rise velocity in turbulence. Since the transition $2.081/Fr_{L}=$ 2.08 already corresponds to $\langle w_{b}\rangle/v_{q}=0.77$ and $\langle w_{b}\rangle/v_{q}$ only reduces further as $1/Fr_{L}$ becomes smaller, using $v_{q}$ as a proxy for $\langle w_{b}\rangle$ (as is done in Berk and Coletti (2024)) would significantly overpredict the value of $\langle w_{b}\rangle$ for most of the tested cases.
@@ 4.2 Standard deviation of the bubble slip velocity component
We next consider the other quantity needed to construct $f(w_{b})$, which is the standard deviation of the slip velocity component $\sigma_{i}=\sqrt{(\sigma_{x}^{2}+\sigma_{y}^{2}+\sigma_{z}^{2})/3}$. Figure 3 shows that $\sigma_{i}$ increases with both $St_{b}$ and $1/Fr$. As $St_{b}$ increases, the bubble is less responsive to higher frequency fluid fluctuation hence $\sigma_{i}$ increases. On the other hand, when $1/Fr$ increases, buoyancy becomes stronger and the bubbles rise faster, which means that they have less time to interact with the same eddy (Csanady, 1963). This ‘crossing trajectory’ effect causes $\sigma_{i}$ to increase. All of the models discussed in §2.2 qualitatively show the same trend in $St_{b}$ as the data, but none of them predict $\sigma_{i}$ increases with $1/Fr$. This is because these models do not account for the crossing trajectory effect, which reduces $T_{L}$ as $1/Fr$ increases (Csanady, 1963). The only reason why $\sigma_{i}^{(BC)}$ (and $\sigma_{i}^{(BCm)}$ which will be introduced in the following paragraph) decreases with increasing $1/Fr$ is due to a larger $f_{b}$ resulting from a larger bubble rise velocity. Although it is possible to include expressions for $T_{L}$ that vary with $1/Fr$ (Berk and Coletti, 2021), it complicates the model as $\sigma_{i}$ would then be different for the vertical and horizontal components. We choose not to include this in our model since it is not of primary importance to the predicted collision kernel in the regime where our model is valid, as will be demonstrated later in this section.
@@
To compare the models quantitatively, we consider the lowest $1/Fr$ case where gravity is weakest as $\sigma_{i}^{(LM)}$ and $\sigma_{i}^{(Kos)}$ do not account for the mean bubble rise velocity. Since the model by Berk and Coletti (2024) depends on the value of the drag correction $f_{b}$, we plot (28) using the measured value of $f_{b}$, i.e. $\sigma_{i}^{(BCm)}$, in addition to $\sigma_{i}^{(BC)}$, where as a first approximation $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ is taken by neglecting the slip velocity fluctuations. Among the models, the one by Berk and Coletti (2024) ($\sigma_{i}^{(BCm)}$) agrees the best with the data across the entire tested range of $St_{b}$, even though in principle it is valid only for $St_{b}\ll 1$; whereas the other models overpredict $\sigma_{i}$ especially when $St_{b}$ is small. For $\sigma_{i}^{(LM)}$, the discrepancy originates from its heuristic approach of balancing drag with buoyancy and the assumption that $\sigma_{i}\sim w_{b}$. As a result, it only gives an order-of-magnitude estimate of $\sigma_{i}$ especially when the correct value of $\zeta$ ($=0.83$) is considered. Meanwhile $\sigma^{(Kos)}$ overpredicts $\sigma_{i}$ because it uses a simpler energy spectrum $E(\omega)$ (see (21)) and neglects the mean bubble rise velocity in the drag correction $f_{b}$. In view of the above, we employ the prediction by Berk and Coletti (2024) in our model of the slip velocity p.d.f.. To keep the model simple, $\sigma^{(BC)}$ instead of $\sigma^{(BCm)}$ will be used. The difference between the two diminishes for large $1/Fr$ as $Re_{b}$ is increasingly dominated by the mean vertical slip velocity.
@@ 4.3 Bubble slip velocity probability density function
Having examined the mean and the standard deviations of the bubble slip velocity components, we focus on the slip velocity magnitude p.d.f. $f(w_{b})$ in figure 4, where the dashed lines show the p.d.f. reconstructed using the measured value of the bubble drag correction $f_{b}$ and will be discussed in §5.1. The model prediction is generally good when $1/Fr\geq 4$ and the small shift in $f(w_{b})$ is due to the slight difference between the actual $\langle w_{b}\rangle$ and the predicted value. This agreement is surprising given our model significantly underpredicts $\sigma_{i}$ when $101/Fr=$ 10 as shown by figure 3. To investigate this, we test our equation for $f(w_{b})$ (13) directly by computing $f(w_{b})$ obtained with the actual $\langle w_{i}\rangle$ and $\sigma_{i}$. The results are shown as dotted curves. These curves agree excellently with the actual $f(w_{b})$ when $1/Fr\leq 2$, but they are wider than the actual $f(w_{b})$ at larger $1/Fr$. This is partly because the value of the vertical component of $\sigma_{i}$ becomes smaller than the horizontal one (Csanady, 1963; Berk and Coletti, 2021), and at higher $1/Fr$, the vertical component dominates the magnitude of the bubble slip velocity, such that the vertical $\sigma_{i}$ contributes more to the shape of $f(w_{b})$. However, the anisotropy of $\sigma_{i}$ does not fully account for the difference from the actual $f(w_{b})$, which suggests that the slip velocity components can only be considered independent when $1/Fr\leq 2$. These limitations are coincidentally compensated by the underprediction of $\sigma_{i}$ in our model, which reduces the width of $f(w_{b})$ and leads to a satisfactory prediction of $f(w_{b})$ at $1/Fr\geq 4$. Our results hence indicate that fully capturing $f(w_{b})$ is far from straightforward and would substantially add to the complexity of the model. The present approach, though relatively simple, gives a reasonable approximation of $f(w_{b})$ and is therefore preferred.
@@
Figure 5: (a) The actual and predicted collision kernels at $St_{b}=2.8$ and (b) $St_{b}=6.3$. Also shown are the collision kernel according to Kostoglou et al. (2020a) $\Gamma^{(Kos)}$ and the collision kernel obtained from (11) using the actual p.d.f. $\Gamma^{f(w_{b})}$. The colour scheme in panel (b) follows that of panel (a). (c) The ratio of the actual collision kernel $\Gamma^{sim}$ and $\Gamma^{f(w_{b})}$ to the model prediction at $St_{b}=2.8$ and (d) $St_{b}=6.3$. The colour scheme in panel (d) follows that of panel (c). (e) Results for $\Gamma$ for $St_{p}=0$ and for varying $1/Fr$ relative to the model prediction. In addition to the simulation data (markers), different variants of the Kostoglou model are included as $\Gamma^{(Kos)}$ (full model), $\Gamma^{(Kos)}_{nS}$ (neglecting small-scale turbulence), and $\Gamma^{(Kos)}_{nSnW}$(neglecting small-scale turbulence and collisions in the bubble wake). (f) The radial fluid velocity at collision distance for a stationary bubble exposed to a uniform freestream $u_{\infty}$ such that $Re_{b}^{\infty}=2r_{b}u_{\infty}/\nu=120$ in the simulations compared to the fit used in $\Gamma^{(Kos)}$. $-u_{r}(r_{c})$ is plotted such that positive values imply flow towards the bubble surface.
4.4 Collision kernel

### v2

@@ 2.3 Modelling the bubble slip velocity probability density function
where $\langle\cdot\rangle$ denotes averaging and $\sigma_{i}$ is the standard deviation of the bubble slip velocity component. This approach has also been taken by Berk and Coletti (2024). For the horizontal components $\langle w_{b,x}\rangle=\langle w_{b,y}\rangle=0$ as buoyancy acts only along the vertical ($z$) direction and hereafter we take $\langle w_{b,z}\rangle=\langle w_{b}\rangle$ for simplicity. We note that $f(w_{b,i})$ is sometimes better described by a stretched exponential function (Masuk et al., 2021). Nonetheless, as the main difference lies in the tails of the distributions, the impact on the predicted collision kernel $\Gamma$ is minimal (Bragg and Collins, 2014; Saffman and Turner, 1956). Additionally, fitting parameters have to be introduced if a stretched exponential function is employed. We therefore choose to model $f(w_{b,i})$ with a Gaussian distribution. Assuming that the components are independent of each other and $\sigma_{x}=\sigma_{y}=\sigma_{z}=\sigma_{i}$ (Berk and Coletti, 2024), we integrate the joint p.d.f. over a spherical shell $S$ with a radius of $\sqrt{w^{2}_{b,x}+w^{2}_{b,y}+w^{2}_{b,z}}$ to obtain
@@
We note that mathematically evaluating Eq. (11) in the limiting of large $\langle w_{b}\rangle$ leads to undefined values, since $\lim_{x\to\infty}\sinh{(x)}/x\to\infty$ and $\lim_{x\to\infty}\exp{(-x^{2})}\to 0$. This limiting case corresponds to a gravity-dominated regime in which the slip velocity vector is oriented nearly vertically and the velocity fluctuation is negligible. Such a scenario does not occur in the turbulent flow conditions we consider. Nevertheless, when the mean slip velocity is relatively large with respect to the fluctuations, the distribution $f(w_{b})$ converges to a normal distribution centred at $\langle w_{b}\rangle$ (see Appendix C). Therefore, we use
$f(w_{b})$ $=\begin{cases}\,\frac{\sqrt{2}w_{b}}{\sqrt{\pi}\sigma\langle w_{b}\rangle}\exp{\bigg(-\frac{w_{b}^{2}+\langle w_{b}\rangle^{2}}{2\sigma^{2}}\bigg)}\sinh{\bigg(\frac{w_{b}\langle w_{b}\rangle}{\sigma^{2}}\bigg)}&\text{if $\langle w_{b}\rangle/\sigma\leq 16$,}\\ \,\frac{1}{\sqrt{2\pi}\sigma}\exp\bigg(-\frac{(w_{b}-\langle w_{b}\rangle)^{2}}{2\sigma^{2}}\bigg)&\text{otherwise.}\end{cases}$ (12)
@@
2.3.1 Mean vertical slip velocity
@@
[…]
@@
2.3.2 Standard deviation of the slip velocity
@@
where the response function of the bubble to turbulent fluctuations is given by
@@
Note that as the expression of $f_{b}$ used tacitly implies $Re_{b}=2r_{b}\sigma_{i}/\nu$, so (21) should only apply for $1/Fr=0$.
@@
where here we choose $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ (Nguyen and Schulze, 2004) with $\langle w_{b}\rangle$ defined by (14). Note that other expressions of $f_{b}$ may be used as desired.
@@ 3.1 Equations and solvers
[…]
at every time step, where $\Pi$ denotes the entire computational domain, $\langle\cdots\rangle_{\Pi}$ indicates spatial averaging over $\Pi$, and $\Pi||$ is the volume of $\Pi$. Hence, the force exerted by the bubble on the fluid is
@@
In contrast to the bubbles, the size of the mineral particles in the flotation processes are comparable to $\eta$, meaning they can be modelled as point-like particles. Here, we initially consider one-way coupled non-settling particles that can overlap with each other and experience the drag, pressure gradient, and added mass forces, which is consistent with the modelling assumptions in Jiang and Krug (2025). Additionally, complementary simulations incorporating particle gravity are performed to examine the effect of particle settling on the collision kernel. The particle dynamics is therefore governed by (13), except all bubble quantities are replaced by those of the particles, and the drag coefficient is $f_{p}=1+0.15Re_{p}^{0.687}$. The particle equation is solved using a finite-difference scheme where the spatial derivatives are discretised using a second-order central finite-difference method and time marching is performed with a second-order Adams–Bashforth scheme using the same time step size as the fluid solver. Since the terms involving $\mathbf{u}$ need to be evaluated at particle positions, their values are interpolated from the Eulerian grid using a tri-linear scheme.
@@ 3.2 Simulation parameters
The equations described in §3.1 are solved to simulate bubbles with $\rho_{b}/\rho_{f}=1/1000$, $St_{b}=0.5-6.3$ and 1/Fr=1101/Fr=1-10 in homogeneous isotropic turbulence with $Re_{\lambda}=64$. Table 1 lists the parameters of the simulations and the turbulence statistics. We limit ourselves to low bubble volume fractions $\phi$ ($\leq 10$ bubbles) where bubbles rarely approach each other. Nonetheless, in case this happens, we employ a soft collision model with a coefficient of restitution of 0.8 to properly capture the lubrication force at small separations and prevent coalescence (Costa et al., 2015). Our simulations are performed in a triply periodic domain. Hence, in principle it is possible for a rising bubble to encounter its own wake, which would affect the statistics. In the wake, the velocity defect $\delta u_{b}$ decays initially with the downstream distance $z$ as $z^{-1}$ until it reaches the same order of magnitude as the turbulent velocity fluctuations at $z_{0}$, beyond which it decays as $z^{-2}$ (Legendre et al., 2006). We take the velocity defect at the bubble position $\delta u_{b}(0)=\langle v_{b,z}\rangle$, define $z_{0}$ by $\delta u_{b}(z_{0})=u^{\prime}$, and estimate the velocity defect one domain length downstream of the bubble $\delta u_{b}(L_{z})=0.01u^{\prime}\ll u^{\prime}$, which is negligibly small. Additionally, we ran an extra simulation with $(St_{b},1/Fr)=(6,10)$ in a longer domain and find that the statistics do not change significantly. These considerations suggest that the bubble’s motion is not significantly affected by its own wake. To ensure sufficient spatial resolution of the boundary layer on the bubbles, we estimate the thickness of the boundary layer at the front stagnation point as $\delta/(2r_{b})=1.13/Re_{b}^{1/2}$ following Johnson and Patel (1999), and resolve this by at least 7 grid points. A grid convergence study (see D) confirmed that this is sufficient for the present cases. The bubbles are initialised on a primitive cubic lattice with the fluid velocity at their positions and allowed to adapt to the flow for 10 large eddy turnover times before their statistics are collected. The instantaneous fluid velocity at bubble position is determined by averaging the fluid velocity within a distance of $3r_{b}$ from the bubble, which for the case of a bubble subjected to a steady uniform flow recovers a value that reaches 90% of the ground truth (Kidanemariam et al., 2013).
@@ 4.1 Mean vertical bubble slip velocity
We compare the predicted values of the mean vertical bubble slip velocity $\langle w_{b}\rangle$ with the actual values in figure 2(a). The data shows that $\langle w_{b}\rangle$ increases with $St_{b}$ and $1/Fr$, which is nicely captured by our model qualitatively and quantitatively. This justifies our use of (14) to model $\langle w_{b}\rangle$, even though (14) was developed for the bubble rise velocity instead of the slip velocity in turbulence. We additionally review the mechanism that reduces the bubble rise velocity in turbulence for the different cases by showing the values of $1/Fr_{L}$ in figure 2(b). For all the tested parameters except $(St_{b},1/Fr)=(2.8,10)$ and $(6.3,10)$, $1/Fr_{L}$ is less than the transition value of 2.08 that separates the two scalings in (14). This suggests that for the vast majority of the simulated cases, nonlinear drag plays a dominant role in reducing the mean vertical rise velocity in turbulence. Since the transition $1/Fr_{L}=2.08$ already corresponds to $\langle w_{b}\rangle/v_{q}=0.77$ and $\langle w_{b}\rangle/v_{q}$ only reduces further as $1/Fr_{L}$ becomes smaller, using $v_{q}$ as a proxy for $\langle w_{b}\rangle$ (as is done in Berk and Coletti (2024)) would significantly overpredict the value of $\langle w_{b}\rangle$ for most of the tested cases.
@@ 4.2 Standard deviation of the bubble slip velocity component
We next consider the other quantity needed to construct $f(w_{b})$, which is the standard deviation of the slip velocity component $\sigma_{i}=\sqrt{(\sigma_{x}^{2}+\sigma_{y}^{2}+\sigma_{z}^{2})/3}$. Figure 3 shows that $\sigma_{i}$ increases with both $St_{b}$ and $1/Fr$. As $St_{b}$ increases, the bubble is less responsive to higher frequency fluid fluctuation hence $\sigma_{i}$ increases. On the other hand, when $1/Fr$ increases, buoyancy becomes stronger and the bubbles rise faster, which means that they have less time to interact with the same eddy (Csanady, 1963). This ‘crossing trajectory’ effect causes $\sigma_{i}$ to increase. All of the models discussed in §2.3.2 qualitatively show the same trend in $St_{b}$ as the data, but none of them predict $\sigma_{i}$ increases with $1/Fr$. This is because these models do not account for the crossing trajectory effect, which reduces $T_{L}$ as $1/Fr$ increases (Csanady, 1963). The only reason why $\sigma_{i}^{(BC)}$ (and $\sigma_{i}^{(BCm)}$ which will be introduced in the following paragraph) decreases with increasing $1/Fr$ is due to a larger $f_{b}$ resulting from a larger bubble rise velocity. Although it is possible to include expressions for $T_{L}$ that vary with $1/Fr$ (Berk and Coletti, 2021), it complicates the model as $\sigma_{i}$ would then be different for the vertical and horizontal components. We choose not to include this in our model since it is not of primary importance to the predicted collision kernel in the regime where our model is valid, as will be demonstrated later in this section.
@@
To compare the models quantitatively, we consider the lowest $1/Fr$ case where gravity is weakest as $\sigma_{i}^{(LM)}$ and $\sigma_{i}^{(Kos)}$ do not account for the mean bubble rise velocity. Since the model by Berk and Coletti (2024) depends on the value of the drag correction $f_{b}$, we plot (26) using the measured value of $f_{b}$, i.e. $\sigma_{i}^{(BCm)}$, in addition to $\sigma_{i}^{(BC)}$, where as a first approximation $f_{b}=1+0.169(2r_{b}\langle w_{b}\rangle/\nu)^{2/3}$ is taken by neglecting the slip velocity fluctuations. Among the models, the one by Berk and Coletti (2024) ($\sigma_{i}^{(BCm)}$) agrees the best with the data across the entire tested range of $St_{b}$, even though in principle it is valid only for $St_{b}\ll 1$; whereas the other models overpredict $\sigma_{i}$ especially when $St_{b}$ is small. For $\sigma_{i}^{(LM)}$, the discrepancy originates from its heuristic approach of balancing drag with buoyancy and the assumption that $\sigma_{i}\sim w_{b}$. As a result, it only gives an order-of-magnitude estimate of $\sigma_{i}$ especially when the correct value of $\zeta$ (=0.83=0.83) is considered. Meanwhile $\sigma^{(Kos)}$ overpredicts $\sigma_{i}$ because it uses a simpler energy spectrum $E(\omega)$ (see (19)) and neglects the mean bubble rise velocity in the drag correction $f_{b}$. In view of the above, we employ the prediction by Berk and Coletti (2024) in our model of the slip velocity p.d.f.. To keep the model simple, $\sigma^{(BC)}$ instead of $\sigma^{(BCm)}$ will be used. The difference between the two diminishes for large $1/Fr$ as $Re_{b}$ is increasingly dominated by the mean vertical slip velocity.
@@ 4.3 Bubble slip velocity probability density function
Having examined the mean and the standard deviations of the bubble slip velocity components, we focus on the slip velocity magnitude p.d.f. $f(w_{b})$ in figure 4, where the coloured dashed lines show the p.d.f. reconstructed using the measured value of the bubble drag correction $f_{b}$ and will be discussed in §5.1. The model prediction is generally good when $1/Fr\geq 4$ and the small shift in $f(w_{b})$ is due to the slight difference between the actual $\langle w_{b}\rangle$ and the predicted value. This agreement is surprising given our model significantly underpredicts $\sigma_{i}$ when $1/Fr=10$ as shown by figure 3.
@@
To investigate this, we test equation (11) directly by computing $f(w_{b})$ obtained with the actual $\langle w_{i}\rangle$ and $\sigma_{i}=\sqrt{(\sigma_{x}^{2}+\sigma_{y}^{2}+\sigma_{z}^{2})/3}$. The results, $f(w_{b})^{(m)}$, are shown as blue dashed curves in figure 4. These curves agree excellently with the actual $f(w_{b})$ when $1/Fr\leq 2$, but they are wider than the actual $f(w_{b})$ at larger $1/Fr$. To pinpoint the source for this discrepancy, we relax the assumption that $\sigma_{x}=\sigma_{y}=\sigma_{z}$ and reconstruct $f(w_{b})$ using the measured values of $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ to obtain $f(w_{b})^{(m)}_{an}$. These are plotted as grey dotted curves for the $1/Fr>2$ cases. The quantities $f(w_{b})^{(m)}_{an}$ and $f(w_{b})^{(m)}$ are almost indistinguishable for all tested cases of $1/Fr\leq 4$ and for cases with $1/Fr=10$ if $St_{b}\leq 1$. This shows that it is appropriate to take $\sigma_{x}=\sigma_{y}=\sigma_{z}$ in this parameter range. Outside this range ($1/Fr<4$), the value of the vertical component ($\sigma_{z}$) becomes noticeably smaller than the horizontal ones (Csanady, 1963; Berk and Coletti, 2021). Additionally, at higher $1/Fr$, the vertical component dominates the magnitude of the bubble slip velocity, such that $\sigma_{z}$ contributes more to the shape of $f(w_{b})$. However, for all the cases where $1/Fr>2$, the anisotropy of $\sigma_{i}$ does not fully account for the difference from the actual $f(w_{b})$. This suggests that the slip velocity components can only be considered independent when $1/Fr\leq 2$. These limitations are coincidentally compensated by the underprediction of $\sigma_{i}$ in our model, which reduces the width of $f(w_{b})$ and leads to a satisfactory prediction of $f(w_{b})$ at $1/Fr\geq 4$.
Our results hence indicate that fully capturing $f(w_{b})$ is far from straightforward and would substantially add to the complexity of the model. The present approach, though relatively simple, gives a reasonable approximation of $f(w_{b})$ and is therefore preferred.
@@
Figure 5: (a) The actual and predicted collision kernels at $St_{b}=2.8$ and (b) $St_{b}=6.3$. Also shown are the collision kernel according to Kostoglou et al. (2020a) $\Gamma^{(Kos)}$ and the collision kernel obtained from (3) using the actual p.d.f. $\Gamma^{f(w_{b})}$. The colour scheme in panel (b) follows that of panel (a). (c) The ratio of the actual collision kernel $\Gamma^{sim}$ and $\Gamma^{f(w_{b})}$ to the model prediction at $St_{b}=2.8$ and (d) $St_{b}=6.3$. The colour scheme in panel (d) follows that of panel (c). (e) Results for $\Gamma$ for $St_{p}=0$ and for varying $1/Fr$ relative to the model prediction. In addition to the simulation data (markers), different variants of the Kostoglou model are included as $\Gamma^{(Kos)}$ (full model), $\Gamma^{(Kos)}_{nS}$ (neglecting small-scale turbulence), and $\Gamma^{(Kos)}_{nSnW}$(neglecting small-scale turbulence and collisions in the bubble wake). (f) The radial fluid velocity at collision distance for a stationary bubble exposed to a uniform freestream $u_{\infty}$ such that $Re_{b}^{\infty}=2r_{b}u_{\infty}/\nu=120$ in the simulations compared to the fit used in $\Gamma^{(Kos)}$. $u_{r}(r_{c})$ is plotted such that positive values imply flow towards the bubble surface.
4.4 Collision kernel with non-settling particles